### v1

# Emergent Mixture-of-Experts: Can Dense Pre-trained Transformers Benefit from Emergent Modular Structures?

Abstract
Incorporating modular designs into neural networks demonstrates superior out-of-generalization, learning efficiency, etc. Existing modular neural networks are generally explicit because their modular architectures are pre-defined, and individual modules are expected to implement distinct functions. Conversely, recent works reveal that there exist implicit modular structures in standard pre-trained transformers, namely Emergent Modularity. They indicate that such modular structures exhibit during the early pre-training phase and are totally spontaneous. However, most transformers are still treated as monolithic models with their modular natures underutilized. Therefore, given the excellent properties of explicit modular architecture, we explore whether and how dense pre-trained transformers can benefit from emergent modular structures. To study this question, we construct Emergent Mixture-of-Experts (EMoE). Without introducing additional parameters, EMoE can be seen as the modular counterpart of the original model and can be effortlessly incorporated into downstream tuning. Extensive experiments (we tune 1785 models) on various downstream tasks (vision and language) and models (22M to1.5B) demonstrate that EMoE effectively boosts in-domain and out-of-domain generalization abilities. Further analysis and ablation study suggest that EMoE mitigates negative knowledge transfer and is robust to various configurations. Code is available at https://github.com/qiuzh20/EMoE
1 Introduction
Modularity attracts considerable attention from the artificial intelligence community (Auda & Kamel, 1999). Neural networks with modular designs, termed Modular Neural Networks (MNN), have exhibited a wide range of advantages, including adaptation (Shen et al., 2023b), out-of-distribution (OOD) generalization abilities (Goyal & Bengio, 2020; Weiss et al., 2022), and data efficiency (Bengio et al., 2020). Typical MNNs are usually explicitly modular. Their modular structure is pre-defined and they are expected to achieve a divide-and-conquer solution for the given task. For example, Andreas et al. (2016) design separate visual and language functional modules and jointly train them to solve visual question-answering tasks. Among various MNNs, Mixture-of-Experts (MoEs) employ a conditional computation strategy where different submodules - so-called experts - are expected to be activated by different types of inputs. MoEs sees substantial success in various domains (Shen et al., 2023a; Chen et al., 2023b; Mustafa et al., 2022; Bao et al., 2022) in the era of large-scale transformers, and therefore becomes a widespread modular neural architecture.
Apart from developing explicit MNN,  Csordás et al. (2021); Agarwala et al. (2021) study whether standard neural networks become implicitly modular after training and discover spontaneously emerged modular structure in small-scale CNNs and LSTMs. For more complicated large-scale pre-trained transformers, initial observations (Zhang et al., 2022b; Li et al., 2022) reveal notable sparse activation patterns within the Feed-Forward Networks (FFNs). Specifically, they find that in T5-Base (Raffel et al., 2020) and ViT-B16, only 3.0% and 6.3% neurons are activated during one forward process. However, sparsity does not imply modularity. Sparsely activated FFNs neurons do not ensure to have remarkable functional division. Therefore,  Zhang et al. (2023) further utilize handpicked semantic and knowledge-intensive tasks to probe the nature of neurons in FFNs. They observe a strong correlation between activation and specific tasks, further discovering clear function-based neuron grouping of the model and summarizing such phenomenon as Emergent Modularity.
Figure 1: (a) Existing literatures (Geva et al., 2021; 2022) suggest that the FFNs in transformers can be viewed as key-value memories. They regarded the input as a query, the first layer as keys, and the second as values. Given an input, keys are sparsely activated (marked in red). Most of the values don’t impact the output. (2) The FFNs block can be partitioned into experts by clustering keys. (3) Afterward, experts’ keys averages are used as the gating weights. The inner product between x and gating weights are used to select experts.
However, contemporary transformers are still employed as monolithic models, with their emergent modularity underutilized. Motivated by the advantages of explicit modularity, we explore whether and how standard pre-trained transformers could benefit from emergent modular structures. Technically, we externalize the emergent modularity into the explicit MoEs, i.e., Emergent Mixcture-of-Experts (EMoE). As shown in Figure 1, we first (1) utilize a clustering-based method to decompose the FFNs into experts and then (2) use the Average-Key-Gating ($\operatorname{avg-k}$ gating) to route the input to corresponding experts. Both two steps do not require extra training processes and data.
Then we fine-tune EMoE on a wide range of downstream tasks to investigate if standard pre-trained transformers can benefit from its ermgent modular structure. We evaluate EMoE following different configurations: (1) tasks from various modalities: vision (Domainbed (Gulrajani & Lopez-Paz, 2021)) and language (GLUE (Wang et al., 2019) and GLUE-X (Yang et al., 2023)); (2) different pre-trained dense transformers: ViT (Dosovitskiy et al., 2021), BERT (Devlin et al., 2019), and GPT2 (Radford et al., 2019), ranging from 22M to 1.5B parameters. (3) different evaluation settings: in-domain (ID) evaluation and OOD evaluation; (4) different fine-tuning methods: full fine-tuning and parameter-efficient tuning (Hu et al., 2022). In total, we fine-tune 1785 models.
Refer to caption Refer to caption
Figure 2: EMoE achieves stronger ID and OOD performances compared with baselines in both full-fine-tuning (Left, average accuracy) and LoRA tuning settings (Hu et al., 2022) (Right, average accuracy for ID and Friedman rank (Friedman, 1940) for OOD). Results are summarized from Section 4.
Our results are summarized in Figure 2. We find that fine-tuning EMoE achieves significant improvement compared with the vanilla fine-tuning and competitive performance with strong baseline GMoE (Li et al., 2023), demonstrating that dense pre-trained transformer can indeed benefit from emergent modular structures to achieve better generalization performance. We further find that EMoE can deactivate neurons with negative transfer effects, thus improving downstream performance. Meanwhile, our ablation studies indicate that EMoE’s insensitivity to various hyper-parameter configurations and higher data efficiency of up to 20%.
2 Related Work
Modular neural networks has various advantages, including interpretability (Pfeiffer et al., 2023), scalability (Chowdhery et al., 2022), multi-task learning abilities (Chen et al., 2023a), and OOD generalization abilities (Goyal et al., 2021; Li et al., 2023). MoEs (Szymanski & Lemmon, 1993) is currently regarded as a standard framework for developing modular neural networks (Shen et al., 2023b). Most MoEs models are developed by training explicit MoEs architecture from random intialization (Shazeer et al., 2017; Fedus et al., 2022b; Zhang et al., 2022a), which usually causes load imbalance issues (Zoph et al., 2022; Fedus et al., 2022a; Chi et al., 2022). Some recent works employ already-trained dense models as a good initialization (Nie et al., 2021). For example, GMoE (Li et al., 2023) and Upcycling Komatsuzaki et al. (2023) copy the FFNs from a trained dense network to continue pre-training or transfer learning. EMoE is also based on pre-trained dense models. However, unlike existing methods that often significantly increase in parameters (GMoE increases 50% parameters) or tend to be less effective (Upcycling extends T-5 Base (248M) to 2B parameters but still underperforms the T5-Large (783M)), EMoE achieves significant improvements without introducing extra training processes or parameters. This is primarily attributed to the heterogeneity of its experts and gating mechanisms, as we demonstrate in Section  5 and 6.
In another research line,  Hod et al. (2021); Csordás et al. (2021) explore modular structures in CNNs and LSTMs. Some works (Zhang et al., 2022b; Li et al., 2022) focus on the sparsity of more complicated dense pre-trained transformers. Based on their observations,  Zhang et al. (2023) recently explore modularity in pre-trained transformer FFNs. In this work, we adopt clustering-based expert construction and $\operatorname{avg-k}$ gating from MoEfication (Zhang et al., 2022b), while our motivation differs. Seeking to preserve the original functionality with fewer activated neurons, they split every layer in the dense model into MoEs. However, the results are worse than the dense model by more than ten points after splitting (Table 2 in their original paper). This is because the clustering-based decomposition and $\operatorname{avg-k}$ gating cannot precisely select the activated neurons. Without any tuning, these errors accumulate with MoEs layers and lead to severe performance dropping. Therefore, Zhang et al. (2022b) mainly focus on more complicated learning-based construction methods. Differently, we treat these decomposed units as heterogeneous modules to exploit their modular properties in the framework of MoEs, and observe significant improvements over monolithic models.
3 Methodology
3.1 Preliminaries
Transformer Feed-Forward Networks The FFNs layer in the transformer block typically includes weights $\mathbf{K}\in\mathbb{R}^{d\times h}$, $\mathbf{V}\in\mathbb{R}^{h\times d}$, and a non-linear activation function $\sigma(\cdot)$. For an input $\mathbf{x}\in\mathbb{R}^{h}$, the output $\mathbf{y}\in\mathbb{R}^{h}$ can be calculated as follows:
$\mathbf{y}=\operatorname{FFN}(\mathbf{x};\mathbf{K},\mathbf{V})=\sigma(\mathbf{x}\cdot\mathbf{K}^{T})\cdot\mathbf{V}.$ (1)
More precisely, for each row $\mathbf{K}_{i,:}$ and column $\mathbf{V_{:,i}}$, the Equation 1 can be rewritten as:
$\mathbf{y}=\sigma(\mathbf{x}\cdot\mathbf{K}^{T})\cdot\mathbf{V}=\sum_{i=1}^{h}\sigma(\mathbf{x}\cdot\mathbf{K_{i,:}})\cdot\mathbf{V_{:,i}}$ (2)
Regarding each row in $\mathbf{K}$ as a key vector and each column in $\mathbf{V}$ as the value vector, the output of an FFNs network can be viewed as a weighted sum of values based on the activation scores $\sigma(\mathbf{x}\cdot\mathbf{K}^{T})$.
Mixture-of-Experts In transformers, MoEs is often implemented by replacing the original FFNs with a group of parallel FFNs and introducing a gating module. Supposing there are $N$ experts: $\{\operatorname{FFN}_{n}(\cdot;\mathbf{K}^{n},\mathbf{V}^{n})|\,n\in[1,N]\}$, the gating module $g(\cdot\,;\mathbf{G},k)$, defined with its parameters $\mathbf{G}$ and an integer $k$, is to map input $\mathbf{x}$ to a score distribution of experts $g(\mathbf{x};\,\mathbf{G},k)\in\mathcal{R}^{N}$. Typically, $g$ is implemented with a simple linear layer followed by a $\operatorname{softmax}$ function and a $\operatorname{Top-k}$ function. Given an input $\mathbf{x}\in\mathbb{R}^{h}$, let $\mathbf{y}_{n}=\operatorname{FFN}_{n}(\mathbf{x};\mathbf{K}^{n},\mathbf{V}^{n})$ be the output of the $n$-th expert, and then the output $\mathbf{y}\in\mathbb{R}^{h}$ of can also be summarized as the weighted sum of the output from all experts:
$\mathbf{y}=\sum_{n\in N}g_{n}(\mathbf{x};\mathbf{G},k)\operatorname{FFN}_{n}(\mathbf{x};\mathbf{K}^{n},\mathbf{V}^{n})=\sum_{n\in N}g_{n}(\mathbf{x};\mathbf{G},k)y_{n},$ (3)
When $k$ for $\operatorname{Top-K}$ is smaller than $N$, only a subgroup of experts is involved in the computation.
3.2 Emergent Mixture-Of-Expert
The ideal EMoE should reflect the implicit modular nature of dense pre-trained transformers using the form of MoEs. A preferred approach to constructing EMoE should avoid introducing additional parameters, training, and data, which may result in undesired biases and are impractical. Although more sophisticated constructing methods can be proposed, we adopt a method from MoEfication Zhang et al. (2022b). The method contains two steps: (1) clustering-based expert construction and (2) $\operatorname{avg-k}$ gating. It is simple and fulfills our needs, as demonstrated in our experiments.
Clustering-based Experts Construction. Since key-value pairs with similar key vectors tend to be co-activated, we split them into separate experts according to their key vectors. Specifically, given a trained FFNs layer $\operatorname{FFNs}(\cdot;\mathbf{K},\mathbf{V})$, we perform constrained clustering (Malinen & Fränti, 2014) to partition all key vectors $\mathbf{K}$ into $N$ experts on average, so each group contains $\frac{d}{N}$ key-value pairs (usually $d=4h$). Denoting the indices of keys in the $i$-th group as $E_{i}\subset[d]$, for $\forall j\in E_{i}$, we extract key-value pair $(\mathbf{K_{j,:}},\mathbf{V_{:,j}})$ to form the $i$th expert $\operatorname{FFN}(\cdot;\mathbf{K}_{i},\mathbf{V}_{i})$, as depicted in Figure 1(b). After that, the computation of each expert proceeds as Equation 1.
Avg-k Gating. Given an input $\mathbf{x}$ and $N$ experts, a qualified gate for EMoE should route the input $\mathbf{x}$ to the experts who contribute most to the model’s output. We construct the gating module by averaging each expert’s keys. The gating function is usually implemented by a single layer $\mathbf{G}\in\mathcal{R}^{h\times N}$, in the $\operatorname{avg-k}$ gating’s case, the weights in $n$th column $\mathbf{G}_{,:n}$ can be calculated as follows:
[…]
And then the corresponding gating score for $i$-th expert is:
$g_{i}(\mathbf{x};\mathbf{G},k)=\begin{cases}&1~{}~{}\text{ if }i\in\operatorname{Top-K}(\mathbf{x}\cdot\mathbf{G}^{T};k,\text{dim=1})\\ &0~{}~{}\text{ else}\end{cases}$ (5)
where $\operatorname{Top-K}(\cdot;k,\text{dim})$ return indices of $k$ largest element of the given input along a given dimension. Using 0,1 score, $\operatorname{avg-k}$ gating reduces the weights’ influence in Equation 3). Notably, observe:
$\mathbf{x}\cdot\mathbf{G}_{:,n}^{T}=\mathbf{x}\cdot\operatorname{Avg}(\mathbf{K}^{n},\text{dim}=0)=\mathbf{x}\cdot\frac{N}{h}\sum_{j\in E_{i}}\mathbf{K}^{n}_{j,:}=\frac{N}{h}\sum_{j}a_{j}.$ (6)
A larger value of gating score $g_{i}$ implies more activated keys within the corresponding expert. Consequently, the expert could potentially contribute more to the output $\mathbf{y}$ for input $\mathbf{x}$. During downstream tuning, gating weights are updated along with the FFNs parameters using Equation 4.
3.3 Discussion
Though adopting a method from MoEfication (Zhang et al., 2022b), EMoE has a distinct motivation, leading to different model configurations and testing scenarios. MoEfication applies the above method to each layer of fine-tuned models, hoping to boost inference efficiency by discarding inactivated neurons. However, the method introduces deviations from FFNs as it can mistakenly eliminate useful neurons. As these deviations accumulated across layers, the performances severely drop by over 10 points on original tasks. Therefore, they focus on more precise expert construction and gating to preserve performances. In sharp contrast, we treat the decomposed experts as heterogeneous modules and explore the benefits of such modularity in downstream tasks. We summarize their main differences in Appendix A.1. As one of our most related work, GMoE (Li et al., 2023) introduce MoEs by copying the trained FFNs. They further demonstrate introducing MoEs every 2 layers leads to significant degradation. Accordingly, we only introduce EMoEs in certain layers since efficiency isn’t our primary concern. We also conduct ablation studies about MoEfying every 2 layers and get consistent results, more details are in Appendix A.4.2 Table 14.
4 Experiments
[…]
Baselines On Domainbed, our baselines include (1) vanilla ViT, as it is a strong OOD baseline under fair configurations suggested by Gulrajani & Lopez-Paz (2021). (2) GMoE (Li et al., 2023), which is the latest state-of-the-art. GMoE constructs MoEs by replicating FFNs from the two-to-last and fourth-to-last transformer blocks of a pre-trained ViT and therefore is the most relevant to EMoE. For language tasks, besides vanilla backbone and GMoE, we implement (3) noisy tuning (Wu et al., 2022), which also improves adaptation for free by adding uniform distribution noise proportional to the standard deviation of the pre-trained weights before tuning. (4) EMoE-learn, an ablation method, where the gating function is learned (same as GMoE) during fine-tuning instead of employing $\operatorname{avg-k}$ gating. This helps us better understand the effect of $\operatorname{avg-k}$ gating.
Table 1: Overall OOD performances with 3 selection criteria. All the reported results are obtained following the Domainbed codebase. The best result is highlighted in bold. In cases where results are the same, the best result is determined by the smallest standard deviation. EMoE demonstrates comparable results to GMoE.
Results with ViT-small (22M) backbone
Results with ViT-base (86M) backbone
Algorithm PACS VLCS
Office
Terra
Avg
Train-validation selection criterion
ViT 86.9 79.7 73.0 44.0 70.90
GMoE 87.7 79.6 73.1 45.4 71.45
EMoE-learn
87.2 79.6 72.5 46.1 71.35
EMoE 87.8 79.5 73.1 45.9 71.58
Leave-one-domain-out selection criterion
ViT 86.1 79.7 73.3 45.0 71.03
GMoE 86.5 80.5 73.1 45.3 71.35
EMoE-learn
86.8 79.6 72.6 45.8 71.20
EMoE 86.8 80.6 73.3 46.1 71.70
Test-domain selection criterion
ViT 86.5 78.2 73.1 44.0 70.45
GMoE 87.2 79.0 73.4 45.3 71.23
EMoE-learn
87.4 79.1 72.8 45.4 71.18
EMoE 87.6 79.0 73.3 45.5 71.35
Algorithm PACS VLCS
Office
Terra
Avg
Train-validation selection criterion
ViT 89.1 80.7 77.2 47.3 73.58
GMoE 90.0 80.4 77.0 49.2 74.15
EMoE-learn
89.8 80.6 76.5 48.7 73.90
EMoE 89.4 80.7 77.3 48.5 73.98
Leave-one-domain-out selection criterion
ViT 88.9 80.8 77.5 46.1 73.33
GMoE 89.3 81.0 76.7 50.1 74.28
EMoE-learn
89.3 81.2 76.5 50.5 74.38
EMoE 89.6 81.6 77.4 50.0 74.65
Test-domain selection criterion
ViT 88.8 79.0 77.2 46.7 72.93
GMoE 89.7 79.0 77.0 48.8 73.63
EMoE-learn
89.7 79.7 76.6 48.7 73.68
EMoE 89.7 79.7 77.5 48.8 73.93
Results with ViT-base (86M) backbone
4.1 Full Fine-tuning Performance
We test the EMoE’s OOD performance on Domainbed. Domainbed provides comprehensive vision OOD evaluations (one result is aggregated with 30 experiments), and the outcomes vary marginally. Moreover,  Gulrajani & Lopez-Paz (2021) indicates that vanilla full fine-tuning with fair hyper-parameter search is a strong baseline compared with specifically designed methods like Invariant Risk Minimization (Arjovsky et al., 2019). More dataset details are in appendix A.2.1. According to Table 1: (1) Overall, EMoE outperforms ViT and GMoE (except ViT-base Train-validation setting, upper right). (2) Compared with EMoE, EMoE-learn incorporates a learned gate. While it surpasses $\operatorname{avg-k}$ gating in certain scenarios (ViT-small Terra Train-validation), it can also lead to a performance drop compared with vanilla ViT. Its overall performance is lower than EMoE. (3) In tasks where the dense model is strong (like Office), EMoE performs better than other MoEs methods. One possible reason is that the $\operatorname{avg-k}$ gating reduces the influence of gating weights ($g_{n}(\mathbf{x};\mathbf{G};k)$ in Equation 3), making it more like the dense model in such scenarios.
Table 2: ID performance on GLUE tasks with different backbones and algorithms. All the reported results are obtained from 3 independent experiments. The average accuracy (Avg) is reported along with the relative improvement compared to the baseline. The best result is highlighted in bold.
Backbone
Algorithm
MRPC CoLA RTE STSB SST2 Avg
BERT Base baseline 88.45 60.67 68.95 87.87 91.97 79.582
noisy tuning
88.43 61.79 71.36 88.27 92.32 80.43(+0.85)
GMoE 88.63 61.25 70.28 88.63 92.28 80.21(+0.63)
EMoE-learn
89.05 62.46 70.40 88.47 92.58 80.59(+1.01)
EMoE 89.45 61.55 69.68 88.71 92.89 80.46(+0.87)
BERT Large baseline 89.82 65.41 74.89 89.87 93.50 82.70
noisy tuning
90.42 64.75 73.41 90.05 93.65 82.46(-0.24)
GMoE 91.24 64.90 74.24 90.00 93.58 82.79(+0.09)
EMoE-learn
90.57 65.51 74.72 90.22 93.73 82.95(+0.25)
EMoE 90.74 65.79 76.17 90.31 93.58 83.32(+0.62)
GPT2 Small baseline 84.46 47.07 67.15 86.29 92.13 75.42
noisy tuning
84.15 46.16 67.51 86.09 92.13 75.21(-0.21)
GMoE 85.07 47.77 67.51 86.57 92.35 75.85(+0.43)
EMoE-learn
85.73 47.24 67.99 86.66 92.35 75.99(+0.57)
EMoE 85.40 48.00 68.95 86.64 92.70 76.34(+0.92)
We evaluate EMoE’s ID performance on 5 GLUE tasks. According to Table 2: (1) On average, EMoE and EMoE-learn outperform other baselines. (2) Among the two methods that do not introduce additional parameters, EMoE significantly outperforms noisy tuning. (3) EMoE provides stable improvements over baselines across different settings, demonstrating its generality.
4.2 LoRA-Tuning performance
Table 3: ID and OOD performance of EMoE and baseline models. All the reported results are obtained from 3 independent experiments. OOD Metrics (averaged over 14 OOD tasks, lower is better) provide additional information for out-of-distribution generalization. The best result is highlighted in bold.
Algorithm MRPC CoLA RTE STSB SST2 QNLI QQP MNLI ID-Avg
OOD
BERT-Large (340M Parameters)
LoRA 89.97 63.40 72.92 90.51 93.16 92.20 87.21 85.40 84.35 4.86
Block 89.34 62.10 71.96 90.39 93.35 92.04 88.45 86.20 84.23(-0.12) 4.95
GMoE 89.45 63.80 72.56 90.29 93.85 92.32 87.99 85.92 84.52(+0.18) 4.04
EMoE-block
89.77 63.25 71.60 90.31 93.69 92.09 88.08 86.21 84.38(+0.03) 5.89
EMoE-learn
89.87 64.00 71.36 90.48 93.65 92.40 87.55 85.62 84.37(+0.02) 4.66
EMoE 90.85 65.33 75.21 90.43 93.50 92.23 87.74 85.43 85.09(+0.74) 4.37
GPT2-XL (1.5B Parameters)
LoRA 86.83 60.88 78.70 89.07 95.18 91.84 87.41 86.93 84.61 5.61
Block 86.59 61.18 79.78 89.08 95.45 91.88 87.71 86.95 84.83(+0.22) 5.13
GMoE 87.02 62.81 79.78 89.21 95.41 92.18 89.10 87.17 85.34(+0.73) 4.33
EMoE-block
87.86 62.88 80.05 89.18 95.49 92.10 89.69 86.87 85.52(+0.91) 5.71
EMoE-learn
87.93 61.50 79.90 89.48 95.18 92.33 89.71 87.00 85.38(+0.77) 4.40
EMoE 87.75 62.27 80.02 89.37 95.41 92.10 89.58 87.06 85.45(+0.84) 3.88
With the increasing scale of pre-trained models, parameter-efficient tuning (Houlsby et al., 2019) gets popular. Observing weaker results with LoRA-tuned ViT-Large on Domainbed compared to ViT-Base, we omit LoRA tuning on Domainbed. GLUE-X (Yang et al., 2023) is a recently introduced OOD dataset for the GLUE benchmark. Therefore, we assess EMoE’s ID performance on GLUE and OOD performance on GLUE-X with the standard LoRA tuning. Notably, LoRA weights are only added in attention, with all other weights frozen. Therefore, the only difference between EMoE and its dense counterpart is FFNs activation. For the OOD metric, we follow GLUE-X (Yang et al., 2023) and employ the Friedman rank (Friedman, 1940) $\text{rank}_{f}=\frac{1}{n}\sum_{i=1}^{n}\text{rank}_{i}$. For each method under the same backbone, $\text{rank}_{i}$ is produced based on the best result and the average result for each dataset. For the 13 OOD tasks used in GLUE-X, each method generates 26 $\text{rank}_{i}$ values. The OOD results presented in Table 3 represent the mean of all these $\text{rank}_{i}$ values. The original results of each task can be found in the Appendix A.4.1. For algorithm configurations, since GMoE copies the FFNs of the pre-trained model to form the MoEs, it is meaningless if the MoEs parameters aren’t tuned. Thus, in Table 3, we conduct experiments with LoRA tuning for GMoE and also tune the transformer block where the original FFNs is replaced. For comparison, we also (1) LoRA-tune dense model and fine-tune the transformer block at the corresponding layer (denoted “Block”); (2) LoRA-tune EMoE with the transformed block also updated (denoted “EMoE-block”).
According to Table 3, EMoE continues to enhance the downstream performance in LoRA tuning: (1) EMoE demonstrates significant enhancements compared to LoRA and doesn’t introduce additional trainable parameters or procedures. Notably, EMoE achieves comparable results with GMoE on BERT-large and outperforms it on GPT2-XL. (2) When the blocks are tuned, EMoE also brings improvements (EMoE-block vs. Block) (3) Consistent with the full-finetuning findings, EMoE exhibits higher stability than EMoE-learn and delivers superior overall results. (4) Besides vision, MoEs also improve OOD performance in language tasks (GMoE vs. Block, EMoE vs. LoRA).
5 Analysis
5.1 How does EMoE improve fine-tuning performance?
Refer to caption     Refer to caption
Figure 3: Left: ID and OOD results compared with LoRA for validating EMoE’s training & inference effects. Right: sparse activated training results with different expert selections.
EMoE achieves notable improvements in both ID and OOD scenarios. However, it is non-trivial that simply transforming the pre-trained FFNs into MoEs before fine-tuning can yield such benefits (especially in LoRA tuning where the MoEs part is totally not updated). Therefore, we investigate the reasons behind the enhanced fine-tuning performance. To decouple the impact of updating expert parameters, we focus on LoRA tuning. We use BERT-Large as the backbone model.
EMoE benefits LoRA weights learning instead of influencing inference. EMoE and dense model only differ in FFNs activations; such difference might (1) directly impact outputs during testing and (2) influence the parameter updating during training. In light of this, we propose two variants and compare them with vanilla LoRA tuning and LoRA tuning with EMoE: (a) LoRA2EMoE: splitting a LoRA-tuned dense model into its MoEs counterpart. If the test results differ from the original ones, we can infer that the sparse activation has an impact during testing. (b) EMoE2LoRA: merging the LoRA weights tuned with EMoE back into the dense. If no changes occur, it implies that the primary factor is the parameter updating. According to Figure 3 Left, doing sparse activation during testing has little impact (LoRA2EMoE). But when merging LoRA weights tuned with EMoE back into the dense, the performance remains significantly better than the vanilla LoRA tuning (EMoE2LoRA vs. LoRA) and is almost identical to EMoE. For full results, please refer to Appendix A.4 Table 11.
EMoE masks neurons with negative transfer impacts. The only difference between EMoE and vanilla LoRA tuning is EMoE blocks some activated neurons during training by $\operatorname{Top-k}$ expert selection. Based on these, we hypothesize that EMoE’s effects stem from preventing negative knowledge transfer from blocked neurons. Therefore, we investigate whether there are such negative transfers. Specifically, we study the following expert selection variants: (1) Bottom-k: select $k=16$ experts who get the lowest scores; (2) Not-top-k: select experts who are not among the top-k experts. These variants are evaluated across 6 tasks from GLUE. The averaged outcomes are in Figure 3 Right. Full results can be found in Appendix A.4 Table 12. LoRA tuning results with “Bottom-k” and “Not-top-k” expert selections are worse than vanilla LoRA tuning. One may attribute this drop to the reduced number of employed neurons. To further assess this, we compare “Top-k” with “Bottom-k”, where an equal number of neurons are used. Notably, “Bottom-k” significantly lags behind “Top-k”. This further corroborates that masked neurons have negative transfer effects.
5.2 Ablation studies
Figure 4: Left: Activations of neurons in FFNs a pre-trained ViT (Dosovitskiy et al., 2021). Right: By clustering the keys in the FFNs and rearranging the action values accordingly, activation patterns of neurons emerge. These co-activated patterns correspond to the modular structure within the pre-trained dense model. With appropriate partitioning, we can transform dense networks into modular networks.
Table 4: Clustering-based and random expert constructing results
Top-k Bottom-K
Cluster +0.92 -0.48
Random -0.11 -0.34
Expert Constructing Methods Figure 4 demonstrates that clustering can decompose modular components within the dense model. To provide further evidence that the EMoE’s improvements stem from leveraging modular features rather than just sparse activation, we compare the results of (1) clustering-based expert construction and (2) random construction within the same setting of Section 3. The relative changes in averaged outcomes compared to the dense baseline are shown in Table 4, while full results can be found in the appendix A.4, Table 12. It’s noteworthy that while cluster top-k exhibits a significant improvement over dense, random top-k is conversely worse than dense baseline. This suggests that random construction can negatively impact gating. Moreover, considering selecting weights with negative transfer under bottom-k selections, it’s observed that cluster bottom-k also achieves lower results. In summary, clustering-based methods can externalize implicit modularity within the dense model. Within suitable frameworks like MoEs, such modularity can facilitate downstream tuning.
Figure 5: Results of 3 tasks for different experts splittings and expert select fractions (top-k / number of experts). ‘e=16’ and ‘e=32’ mean splitting FFNs into 16 and 32 experts, respectively.
Expert Constructing Configurations Beyond the settings detailed in the main results section, which are based on $N=64$ experts and top-k $\in\{16,32,48\}$, we also present specific scenarios where $N\in\{16,32\}$, and top-k varies within $\{2,4,8,16\}$ in Figure 5. Notably, in each of these settings, (1) EMoE consistently surpasses the dense model, illustrating its insensitivity to hyper-parameters. (2) On average, $\operatorname{avg-k}$ gating exhibits superior performance than learned gating. Though learned gating (EMoE-learn) outperforms $\operatorname{avg-k}$ gating in a few specific settings (Figure 5 (b) and (e)). This is consistent with the earlier results in Section 4. Regarding how many EMoE layers should be introduced, our findings align with those discovered in GMoE, indicating that only a limited number of layers can be converted into the EMoE layer. If excessive EMoE layers are introduced, performance deteriorates. Taking GPT2-XL (48 layers) as an example, when introducing EMoE every 2 layers in the latter half, the performance averaged across 5 GLUE tasks (79.36) matches that of the dense model (78.87). But when adopting EMoR every 2-layer for the entire model, the performance lags slightly behind that of the dense model (78.17) but surpasses EMoE-learn (75.87). For additional configurations, please refer to the appendix A.4.2 Table 14.
Figure 6: Expert selections during training with distinct gating functions ($\operatorname{avg-k}$ vs. learned) and expert types (modules from dense vs. repetitions of dense). The vertical axis illustrates training steps (top-down arrangement signifies begin-end); the horizontal axis represents expert selection frequency within 1K steps (deeper color implies a higher frequency). (a), (b) and (c) correspond to EMoE, EMoE-learn, and GMoE.
Expert Selection Changes During Training To further understand $\operatorname{avg-k}$ gating and learned gating, we visualize expert selections of GPT2-XL during fine-tuning on 6 tasks with 16 experts. In Figure 6, we showcase the results for the largest dataset QNLI among them. Full results are available in the appendix A.5. Our observations are: (1) Both $\operatorname{avg-k}$ gating and learned gating converge, as indicated by the lower halves of the plots. (2) $\operatorname{avg-k}$ gating is more stable than learned gating (Figure 6 a vs. b). This might mitigate data inefficiency from gating inconsistencies across different stages of training (Zuo et al., 2022). (3) EMoE, with its heterogeneous experts, exhibits better load balancing than GMoE (Figure 6 b vs. c). In GMoE, all experts share identical initialization, whereas in EMoE, the experts are derived from FFNs with implicit modularity. This also suggests a good initialization can facilitate MoEs learning similar to EvoMoE (Nie et al., 2021).
Figure 7: Average performance of EMoE with different proportions of training data.
Performance Across Different Training set Volumes. Previous research has indicated that modular architectures offer improved data efficiency (Bengio et al., 2020). Therefore, we conducted experiments with GPT2-XL on six tasks using varying proportions of original training data, and the results for all tasks are presented in Figure 7. It can be observed that EMoE consistently outperforms the dense across different data factions. EMoE achieves superior results even when using less than 20% of the data. On SST2, only using 50% data, EMoE shows comparable performance to the dense. More details can be found in the Appendix A.4.2 Table 15. This further underscores the benefits of incorporating modular structures.
6 Conclusion, Limitations and Future Works
Conclusion: In this work, we validate that exploiting the emergent modular structures in dense pre-trained transformers improves downstream task ID and OOD performances. One possible reason is the modular structure can alleviate negative transfer effects presented in the pre-trained model. We hope our findings could deepen the understanding of neural networks’ modularity, further helping the community develop more sophisticated modular neural architectures and utilizing existing pre-trained models. Limitations: We would like to point out that EMoE is not designed to outperform the SOTA tuning methods and OOD generalization algorithms. Instead, the goal is to demonstrate that dense pre-trained transformers can benefit from emergent modular structures. Our research findings have not been validated on more challenging tasks (e.g., Mathematical Reasoning (Imani et al., 2023)) or foundation models like LLaMA (Touvron et al., 2023) due to computational constraints. While our evaluation was primarily conducted on models with a maximum parameter count of 1.5B, it is evident that across the spectrum of parameter sizes, ranging from 22M to 1.5B, EMoE consistently yields improvements, showing the potential to scale up. Future works: Subsequent research could delve into better methods for modular decomposition and investigate the modularity in foundation models. Benefiting from EMoE’s ability to avoid negative transfer, future researchers can also exploit the emergent modularity in settings such as multi-task learning and continual learning. Moreover, it is observed that modularity structures emergent and remain stable after approximately one-fourth of the pre-training phase (Li et al., 2022; Zhang et al., 2023). Future investigations could investigate leveraging the modularity of dense models during the pre-training process.
7 Reproducibility
In the Experiment Configuration Section 4, we first introduce the additional configurations introduced by EMoE relative to the Dense Model, as well as the baseline settings. Subsequently, we provide high-level experiment settings for each experimental group (Domainbed at 4.1, full fine-tuning GLUE tasks at section 2, LoRA tuning at  3). In Appendix A.2, we provide detailed information on the datasets used, the codebase, and the evaluation metrics. We also outline more general configurations, such as learning rates and batch sizes for the respective tasks. Beyond the main paper, we include the original code for the experiments and log files for certain experimental results in the supplementary materials. Each section of code includes a README.md file that explains the experimental settings required to replicate the results.
Appendix A Appendix
A.1 Compare with Moefication
Table 5: Comparison between EMoE and Moefication
Aspect EMoE Moefication
Motivation
Modularity
Sparsity
Goal
Transfer learning effective
Inference efficiency
MoEs Layers
Typically 1 or 2 layers
Applies changes across all layers
Scenarios
Down-stream tasks
Original task
Results
Significant improvement without adding parameters
Improved inference efficiency but performance drop
Though adopting the MoEs construction method from MoEfication (Zhang et al., 2022b), EMoE has a distinct motivation, leading to different model configurations and testing scenarios. MoEfication applies the above method to each layer of fine-tuned models, hoping to boost inference efficiency by discarding inactivated neurons. In sharp contrast, we treat the decomposed experts as heterogeneous modules and explore the benefits of such modularity in downstream tasks. Accordingly, We only introduce EMoEs in certain layers according to GMoE, since efficiency is not our primary concern. These differences are summarized in Table 5.
A.2 Datasets and Evaluation Metrics
A.2.1 Domainbed
Table 6: Used dataset information from Domainbed
Dataset PACS VLCS OfficeHome TerraInc
#Domains 4 4 4 4
Classes 7 5 65 10
Images 9,991 10,729 15,588 24,788
The four datasets (PACS (Li et al., 2017), VLCS (Fang et al., 2013), Office-Home (Venkateswara et al., 2017), and Terra Incognita (Beery et al., 2018)) are selected from Domainbed. Each dataset comprises 4 distinct domains(PACS: {art, cartoons, photos, sketches}, VLCS: {Caltech101, LabelMe, SUN09, VOC2007}, Office-Home: {art,clipart, product, real}, and Terra Incognita: {L100, L38, L48, L46}). Within a single training, one or two domains’ data are sequentially designated for OOD evaluation. For example, when training on PACS, {art, cartoons} could be selected as ID training data, while {photos, sketches} are designated for OOD testing. This configuration results in a total of $C_{4}^{2}+C_{4}^{1}=10$ distinct training processes within each dataset. Suppose there are $d_{tr}$ ID domains, “Train-validation” means selecting OOD test checkpoints based on ID accuracies from the validation subsets of all $d_{tr}$ ID domains; “Leave-one-domain-out” means leaving one selected ID domain as a validation set, doing training on $d_{tr}-1$ domains; “Test-domain” means selection based on limited accesses to test domains and selecting based on these results. The final results are aggregated with selection criteria provided by Domainbed [^1] https://github.com/facebookresearch/DomainBed. As a result, even a variation of 0.1 in the benchmark outcomes signifies a significant improvement.
In our experiments, all the hyper-parameters, like training steps, learning rates, and weight decay, except those related to MoEs, strictly follow GMoE.
A.2.2 GLUE
Each task involves one to four OOD tasks from GLUE-X (Yang et al., 2023), resulting in 13 OOD tasks in total. To illustrate, consider the Sentiment Analysis task: we first fine-tune models on SST-2 from GLUE and report the validation results as ID performance, then use the test data from IMDB (Maas et al., 2011), Yelp (Zhang et al., 2015), Amazon (Kaushik et al., 2020) and Flipkart (Vaghani & Thummar, 2023) from GLUE-X for OOD testing.
In the full fine-tuning, to ensure convergence and reduce randomness, we train all models 10 epochs across 3 random seeds on each task. Each experiment does a hyper-parameter search on learning rates on [2e-5, 3e-5, 5e-5] as suggested by BERT (Devlin et al., 2019). The training batch size is 32. In the LoRA tuning, following LoRA (Hu et al., 2022) that tunes models with more epochs and larger learning rates than standard full fine-tuning, all models are trained 20 epochs on small and medium datasets and 5 epochs on large ones (like QNLI, MNLI, QQP). The learning rate is searched in [2e-4, 3e-4, 5e-4]. All methods are implemented with LoRA_rank=8 and LoRA_alpha=16. The training batch size is 16 due to a larger model size. Other settings like max_lengt following the codebase from hugging face [^2] https://github.com/huggingface/transformers/tree/main/examples/pytorch/text-classification. After training on GLUE, we directly test the selected models on GLUE-X with the data from the official repo [^3] https://github.com/YangLinyi/GLUE-X.
Table 7: Language tasks and corresponding ID and OOD datasets.
Task ID-dataset size OOD-dataset size
Paraphrase MRPC 4,076 Twitter 16,777
QQP 404,301 Twitter 16,777
MRPC 4,076
Linguistic Acceptability CoLA 9,594 Grammar Test 304,277
Textual Entailment RTE 2,768 SciTail 26,527
HANs 60,000
Textual Similarity STSB 7,128 SICK 9,840
Sentiment Analysis SST2 68,223 IMDB 50,000
Yelp 598,000
Amazon 4,000,000
Flipkart 205,041
Question Answering NLI QNLI 110,206 NewsQA 119,525
Natural Language Inference MNLI 412,313 SICK 9,840
A.3 Computation Cost and Memory Usage
Theoretically, EMoE does not introduce additional parameters compared to its dense counterpart. Although it adds computation in the gating portion within the MoEs layer, it omits a substantial amount of computation within the FFNs layer. For instance, the computation in the gating portion is of the order of $h\times N$, where h represents the model’s hidden size, and N is the number of experts. In contrast, the complete computation in the FFNs layer is of the order of $(h\times h\times 4h)\times 2$, and sparse activations can reduce more than a quarter of this computation. Since $N\ll h$, theoretically, using EMoE within a single block should accelerate the forward pass of the model. However, in real deployment, we have observed that the hardware implementation of MoEs can result in EMoE being, on average, approximately 10% slower than the dense model. Additionally, the memory usage is also slightly higher, by less than 5%, compared to the dense model (16295MB v.s. 15865MB on CoLA, LoRA tuning GPT2-XL).
Each individual experiment was conducted on a single NVIDIA 40G A100 GPU. The training times for different tasks ranged from just over ten minutes (RTE) to more than ten hours (QQP).
A.4 Tabular Results
A.4.1 Full Tables in Full Fine-tuning with Standard Deviation
In this section, we present the mean and variance of experiments conducted with three different random seeds. The Domainbed results are demonstrated in Table 8, full fine-tuning results are in Table 9, LoRA tuning results are in Table 10.
Table 8: Overall out-of-domain performances with different selection criteria. All the reported results are obtained from three independent experiments conducted following the Domainbed codebase. The best result is highlighted in bold.In cases where results are the same, the best result is determined by the smallest standard deviation. EMoE demonstrates comparable results to GMoE.
Results with ViT-small backbone
Results with ViT-base backbone
Algorithm PACS VLCS OfficeHome TerraInc Avg
train-validation selection criterion
ViT 86.9
$\pm$0.2
79.7 $\pm$0.4 73.0
$\pm$0.2
44.0
$\pm$1.1
70.90
GMoE 87.7
$\pm$0.2
79.6
$\pm$0.4
73.1
$\pm$0.3
45.4
$\pm$0.3
71.45
EMoE-learn
87.2
$\pm$0.4
79.6
$\pm$0.2
72.5
$\pm$0.2
46.1 $\pm$0.4 71.35
EMoE 87.8 $\pm$0.2 79.5
$\pm$0.4
73.1 $\pm$0.2 45.9
$\pm$0.3
71.58
leave-one-domain-out selection criterion
ViT 86.1
$\pm$0.6
79.7
$\pm$0.4
73.3 $\pm$0.1 45.0
$\pm$0.5
71.03
GMoE 86.5
$\pm$0.3
80.5
$\pm$0.2
73.1
$\pm$0.3
45.3
$\pm$0.6
71.35
EMoE-learn
86.8 $\pm$0.0 79.6
$\pm$0.3
72.6
$\pm$0.2
45.8
$\pm$0.6
71.20
EMoE 86.8
$\pm$0.1
80.6 $\pm$0.4 73.3
$\pm$0.2
46.1 $\pm$0.6 71.70
test-domain selection criterion
ViT 86.5
$\pm$0.4
78.2
$\pm$0.4
73.1 $\pm$0.2 44.0
$\pm$0.5
70.45
GMoE 87.2
$\pm$0.4
79.0
$\pm$0.3
73.4 $\pm$0.2 45.3
$\pm$0.4
71.23
EMoE-learn
87.4
$\pm$0.2
79.1 $\pm$0.3 72.8
$\pm$0.1
45.4
$\pm$0.6
71.18
EMoE 87.6 $\pm$0.5 79.0
$\pm$0.2
73.3
$\pm$0.0
45.5 $\pm$0.1 71.35
Algorithm PACS VLCS OfficeHome TerraInc Avg
train-validation selection criterion
ViT 89.1
$\pm$0.0
80.7 $\pm$0.1 77.2
$\pm$0.1
47.3
$\pm$0.8
73.58
GMoE 90.0 $\pm$0.3 80.4
$\pm$0.6
77.0
$\pm$0.1
49.2 $\pm$1.1 74.15
EMoE-learn
89.8
$\pm$0.2
80.6
$\pm$0.2
76.5
$\pm$0.1
48.7
$\pm$0.5
73.90
EMoE 89.4
$\pm$0.4
80.7
$\pm$0.2
77.3 $\pm$0.1 48.5
$\pm$0.5
73.98
leave-one-domain-out selection criterion
ViT 88.9
$\pm$0.4
80.8
$\pm$0.3
77.5 $\pm$0.1 46.1
$\pm$0.6
73.33
GMoE 89.3
$\pm$0.6
81.0
$\pm$0.3
76.7
$\pm$0.1
50.1
$\pm$1.1
74.28
EMoE-learn
89.3
$\pm$0.2
81.2
$\pm$0.1
76.5
$\pm$0.1
50.5 $\pm$0.2 74.38
EMoE 89.6 $\pm$0.2 81.6 $\pm$0.2 77.4
$\pm$0.1
50.0
$\pm$1.1
74.65
test-domain selection criterion
ViT 88.8
$\pm$0.7
79.0
$\pm$0.5
77.2
$\pm$0.0
46.7
$\pm$0.4
72.93
GMoE 89.7
$\pm$0.5
79.0
$\pm$0.3
77.0
$\pm$0.1
48.8 $\pm$0.4 73.63
EMoE-learn
89.7 $\pm$0.4 79.7 $\pm$0.2 76.6
$\pm$0.1
48.7
$\pm$0.3
73.68
EMoE 89.7 $\pm$0.4 79.7 $\pm$0.2 77.5 $\pm$0.1 48.8
$\pm$0.6
73.93
Results with ViT-base backbone
Table 9: Results on GLUE tasks with different backbones and algorithms. All the reported results are obtained from 3 independent experiments. The average accuracy (avg) is reported along with the relative improvement compared to the baseline. The best result is highlighted in bold.
Backbone
Algorithm
MRPC CoLA RTE STSB SST2 Avg
BERT-B baseline 88.45
$\pm$0.40
60.67
$\pm$0.54
68.95
$\pm$0.69
87.87
$\pm$0.12
91.97
$\pm$0.19
79.582
noisy tuning
88.43
$\pm$0.12
61.79
$\pm$0.16
71.36
$\pm$0.17
88.27
$\pm$0.94
92.32
$\pm$0.25
80.43(+0.85)
GMoE 88.63
$\pm$0.53
61.25
$\pm$2.36
70.28
$\pm$0.68
88.63
$\pm$0.65
92.28
$\pm$0.24
80.21(+0.63)
EMoE-learn
89.05
$\pm$0.23
62.46 $\pm$1.01 70.40 $\pm$1.28 88.47
$\pm$0.74
92.58
$\pm$0.14
80.59(+1.01)
EMoE 89.45 $\pm$0.36 61.55
$\pm$0.67
69.68
$\pm$1.02
88.71 $\pm$0.50 92.89 $\pm$0.19 80.46(+0.87)
BERT-L baseline 89.82
$\pm$1.30
65.41
$\pm$0.47
74.89
$\pm$1.39
89.87
$\pm$0.28
93.50
$\pm$0.24
82.70
noisy tuning
90.42
$\pm$0.35
64.75
$\pm$1.31
73.41
$\pm$1.62
90.05
$\pm$0.46
93.65
$\pm$0.11
82.46(-0.24)
GMoE 91.24 $\pm$0.25 64.90
$\pm$1.26
74.24
$\pm$1.04
90.00
$\pm$0.64
93.58
$\pm$0.25
82.79(+0.09)
EMoE-learn
90.57
$\pm$0.43
65.51
$\pm$0.32
74.72
$\pm$2.13
90.22
$\pm$0.49
93.73 $\pm$0.35 82.95(+0.25)
EMoE 90.74
$\pm$0.65
65.79 $\pm$1.16 76.17 $\pm$0.00 90.31 $\pm$0.43 93.58
$\pm$0.32
83.32(+0.62)
GPT2 baseline 84.46
$\pm$0.51
47.07
$\pm$1.60
67.15
$\pm$0.51
86.29
$\pm$0.29
92.13
$\pm$0.30
75.42
noisy tuning
84.15
$\pm$0.92
46.16
$\pm$2.79
67.51
$\pm$0.78
86.09
$\pm$0.38
92.13
$\pm$0.27
75.21(-0.21)
GMoE 85.07
$\pm$0.45
47.77
$\pm$3.20
67.51
$\pm$0.51
86.57
$\pm$0.29
92.35
$\pm$0.35
75.85(+0.43)
EMoE-learn
85.73 $\pm$0.09 47.24
$\pm$1.48
67.99
$\pm$0.17
86.66 $\pm$0.32 92.35
$\pm$0.38
75.99(+0.57)
EMoE 85.40
$\pm$0.77
48.00 $\pm$1.50 68.95 $\pm$0.29 86.64
$\pm$0.16
92.70 $\pm$0.22 76.34(+0.92)
Table 10: Results on various algorithms with different models and tasks. All the reported results are obtained from 3 independent experiments. OOD Metrics (averaged over 14 OOD tasks, lower is better) provide additional information for out-of-distribution generalization. The best result is highlighted in bold, and the second is marked with underline.
Algorithm MRPC CoLA RTE STSB SST2 QNLI QQP MNLI ID-Avg
OOD
BERT-Large (340 Million Parameters) Results
LoRA 89.97
$\pm$0.40
63.40
$\pm$0.62
72.92
$\pm$1.64
90.51 $\pm$0.18 93.16
$\pm$0.19
92.20
$\pm$0.13
87.21
$\pm$0.60
85.40
$\pm$0.07
84.35 4.86
Block 89.34
$\pm$0.84
62.10
$\pm$0.91
71.96
$\pm$1.68
90.39
$\pm$0.14
93.35
$\pm$0.43
92.04
$\pm$0.16
88.45 $\pm$0.07 86.20 $\pm$0.10 84.23(-0.12) 4.95
Block+GMoE 89.45
$\pm$0.72
63.80
$\pm$0.71
72.56
$\pm$0.29
90.29
$\pm$0.07
93.85 $\pm$0.11 92.32
$\pm$0.14
87.99
$\pm$0.06
85.92
$\pm$0.13
84.52(+0.18) 4.04
Block+EMoE-learn
89.79
$\pm$0.23
64.16
$\pm$0.87
73.16
$\pm$1.04
90.27
$\pm$0.03
93.85 $\pm$0.11 92.40 $\pm$0.06 88.01
$\pm$0.12
85.76
$\pm$0.19
84.68(+0.33) 3.94
Block+EMoE 89.77
$\pm$0.46
63.25
$\pm$0.50
71.60
$\pm$0.68
90.31
$\pm$0.09
93.69
$\pm$0.32
92.09
$\pm$0.13
88.08
$\pm$0.19
86.21 $\pm$0.16 84.38(+0.03) 5.89
EMoE 90.85 $\pm$0.61 65.33 $\pm$0.40 75.21 $\pm$1.62 90.43
$\pm$0.06
93.50
$\pm$0.33
92.23
$\pm$0.10
87.74
$\pm$0.10
85.43
$\pm$0.10
85.09(+0.74) 4.37
EMoE+LN 90.47 $\pm$0.33 64.39 $\pm$0.31 73.41 $\pm$1.04 90.54 $\pm$0.03 93.00
$\pm$0.16
92.31
$\pm$0.05
88.79 $\pm$0.17 85.50
$\pm$0.10
84.80(+0.46) 4.53
EMoE-learn 89.87
$\pm$0.50
64.00
$\pm$0.57
71.36
$\pm$1.39
90.48
$\pm$0.10
93.65
$\pm$0.33
92.40 $\pm$0.11 87.55
$\pm$0.14
85.62
$\pm$0.23
84.37(+0.02) 4.66
EMoE-learn+LN
89.9
$\pm$0.25
64.16
$\pm$1.16
72.44
$\pm$0.45
90.45
$\pm$0.10
93.42
$\pm$0.38
92.15
$\pm$0.10
87.70
$\pm$0.04
85.52
$\pm$0.24
84.47(0.12) 4.28
GPT2-XL (1.5 Billion Parameters) Results
LoRA 86.83
$\pm$0.87
60.88
$\pm$2.54
78.70
$\pm$0.59
89.07
$\pm$0.11
95.18
$\pm$0.28
91.84
$\pm$0.09
87.41
$\pm$1.74
86.93
$\pm$0.15
84.61 5.61
Block 86.59
$\pm$1.45
61.18
$\pm$1.74
79.78
$\pm$2.22
89.08
$\pm$0.15
95.45 $\pm$0.19 91.88
$\pm$0.05
87.71
$\pm$2.95
86.95
$\pm$0.08
84.83(+0.22) 5.13
Block+GMoE 87.02
$\pm$0.76
62.81
$\pm$1.51
79.78
$\pm$1.35
89.21
$\pm$0.20
95.41
$\pm$0.28
92.18
$\pm$0.11
89.10
$\pm$0.78
87.17 $\pm$0.20 85.34(+0.73) 4.33
Block+EMoE-learn
87.31
$\pm$1.23
62.24
$\pm$1.51
79.54
$\pm$0.17
89.33
$\pm$0.11
95.30
$\pm$0.09
92.20
$\pm$0.09
88.59
$\pm$1.68
87.06 $\pm$0.18 85.20(+0.59) 4.05
Block+EMoE 87.86
$\pm$0.98
62.88 $\pm$0.54 80.05 $\pm$0.29 89.18
$\pm$0.25
95.49 $\pm$0.39 92.10
$\pm$0.15
89.69
$\pm$0.15
86.87
$\pm$0.11
85.52(+0.91) 5.71
EMoE 87.75
$\pm$0.14
62.27
$\pm$0.93
80.02 $\pm$0.34 89.37
$\pm$0.30
95.41
$\pm$0.32
92.10
$\pm$0.15
89.58
$\pm$0.10
87.06
$\pm$0.25
85.45(+0.84) 3.88
EMoE+LN 88.05 $\pm$0.35 63.11 $\pm$0.51 79.90
$\pm$1.51
89.40
$\pm$0.22
95.18
$\pm$0.28
92.23
$\pm$0.11
89.70
$\pm$0.09
87.03
$\pm$0.14
85.58(+0.97) 4.39
EMoE-learn 87.93 $\pm$0.61 61.50
$\pm$1.09
79.90
$\pm$0.61
89.48 $\pm$0.24 95.18
$\pm$0.11
92.33 $\pm$0.093 89.71 $\pm$0.06 87.00
$\pm$0.19
85.38(+0.77) 4.40
EMoE-learn+LN
87.04
$\pm$1.11
62.64
$\pm$0.84
79.78
$\pm$0.59
89.50 $\pm$0.22 95.30
$\pm$0.50
92.31 $\pm$0.19 89.43
$\pm$0.35
87.00
$\pm$0.12
85.38(+0.77) 3.67
A.4.2 Full Tables and Figure Data Sources in Analysis
In the Analysis section 5.1, for the sake of clarity, we have transformed tabular data into graphs or retained only a subset of the results. The original and complete results corresponding to them are presented in this section. Figure 3 is summarized from Table 11 and Table 12. The OOD results in Table 3 are from 10. The ablation studies are from Table 14 and Table 15.
Table 11: ID and OOD results of BERT-L for different settings. ”LoRA-to-EMoE” refers to converting a model tuned using standard LoRA into EMoE for testing. On the other hand, ”EMoE-to-LoRA” involves merging a tuned EMoE model back into a standard dense model during testing.
Algorithm CoLA Gram MRPC Twitter RTE Hans SciTail STSB Sick Avg
LoRA 60.89
$\pm$2.55
41.77
$\pm$1.62
86.83
$\pm$0.87
75.42
$\pm$2.71
78.70
$\pm$1.02
60.37
$\pm$1.31
77.36
$\pm$0.73
89.22
$\pm$0.13
78.48
$\pm$0.33
72.12
LoRA-to-EMoE
61.31
$\pm$2.14
41.99
$\pm$1.56
86.83
$\pm$0.91
75.15
$\pm$3.02
78.58
$\pm$0.95
60.39
$\pm$1.32
77.24
$\pm$0.68
89.23
$\pm$0.13
78.53
$\pm$0.35
72.14
EMoE 62.69
$\pm$0.91
42.95
$\pm$0.95
87.82
$\pm$0.17
76.07
$\pm$2.12
79.54
$\pm$0.45
61.56
$\pm$1.65
78.09
$\pm$0.56
89.39
$\pm$0.31
78.57
$\pm$0.67
72.96
EMoE-to-LoRA
62.69
$\pm$0.91
42.94
$\pm$0.95
87.82
$\pm$0.17
76.06
$\pm$2.12
79.54
$\pm$0.45
61.56
$\pm$1.65
78.07
$\pm$0.58
89.39
$\pm$0.3
78.58
$\pm$0.67
72.96
Table 12: ID results of BERT-L for different settings. ”Cluster-top” refers to EMoE utilizing $\operatorname{avg-k}$ gating. ”Cluster-not-top” represents a scenario where, during gating, the top-k experts are removed. Similarly, ”Cluster-bottom” involves selecting the bottom-k experts with the lowest scores during gating. ”Random” denotes the approach of randomly selecting key values to construct experts. The terms ”top,” ”not-top,” and ”bottom” have the same meanings as in the cluster situations.
Algorithm MRPC CoLA RTE STSB SST2 QNLI Avg
LoRA 89.97
$\pm$0.40
63.40
$\pm$0.62
72.92
$\pm$1.64
90.51
$\pm$0.18
93.16
$\pm$0.19
92.20
$\pm$0.13
83.69
Cluster-top 90.85
$\pm$0.61
65.33
$\pm$0.40
75.21
$\pm$1.62
90.54
$\pm$0.03
93.50
$\pm$0.33
92.23
$\pm$0.10
84.61(+0.92)
Cluster-not-top 89.61
$\pm$0.76
63.21
$\pm$0.44
72.56
$\pm$1.28
90.31
$\pm$0.07
93.12
$\pm$0.34
92.14
$\pm$0.18
83.49(-0.20)
Cluster-bottom 89.21
$\pm$0.69
63.08
$\pm$1.09
71.72
$\pm$0.34
90.15
$\pm$0.18
92.97
$\pm$0.19
92.13
$\pm$0.31
83.21(-0.48)
Random-top 89.88
$\pm$0.75
63.26
$\pm$0.39
72.56
$\pm$1.06
90.33
$\pm$0.05
93.35
$\pm$0.00
92.14
$\pm$0.20
83.59(-0.11)
Random-not-top 90.09
$\pm$0.75
63.35
$\pm$0.34
72.44
$\pm$1.19
90.44
$\pm$0.07
93.31
$\pm$0.25
92.20
$\pm$0.16
83.64(-0.05)
Random-bottom 89.47
$\pm$0.23
63.17
$\pm$0.98
71.96
$\pm$0.74
90.30
$\pm$0.23
93.11
$\pm$0.25
92.10
$\pm$0.11
83.35(-0.34)
Table 13: Raw OOD performances across 13 tasks. Average results with standard deviation and best results are reported separately. Due to the large deviation across seeds overall methods, we use Friedman rank metrics (Friedman, 1940).
Algorithm
Twitter-M
GrammarTest
Hans SciTail Sick-S
NewsQA
Amazon Flipkart Imdb Yelp MRPC
Twitter-Q
Sick-M
BERT-Large (340 Million Parameters) Results (mean and standard deviation)
LoRA 80.09
$\pm$0.49
43.55
$\pm$1.34
55.51
$\pm$1.94
81.13
$\pm$0.88
81.63
$\pm$0.16
78.22
$\pm$0.17
89.06
$\pm$1.81
91.79
$\pm$0.27
84.97
$\pm$1.32
88.40
$\pm$2.06
71.73
$\pm$0.46
78.10
$\pm$0.33
53.17
$\pm$0.92
LoRA+block 80.70
$\pm$0.35
42.94
$\pm$0.47
54.35
$\pm$1.08
80.99
$\pm$0.33
81.58
$\pm$0.18
78.33
$\pm$0.24
88.21
$\pm$1.20
91.26
$\pm$0.96
84.10
$\pm$0.85
89.13
$\pm$0.20
71.57
$\pm$0.87
79.22
$\pm$0.59
53.67
$\pm$0.77
GMoE 81.17
$\pm$0.54
46.76
$\pm$1.91
58.73
$\pm$0.74
79.41
$\pm$0.10
81.35
$\pm$0.34
78.05
$\pm$0.28
89.27
$\pm$1.38
91.77
$\pm$0.07
84.93
$\pm$1.09
88.70
$\pm$1.77
71.08
$\pm$0.69
78.70
$\pm$0.37
53.85
$\pm$1.27
Block+EMoE-learn
81.09
$\pm$0.47
43.29
$\pm$1.42
57.55
$\pm$3.48
77.72
$\pm$0.61
80.78
$\pm$0.49
78.32
$\pm$0.13
90.17
$\pm$0.32
91.82
$\pm$0.18
85.71
$\pm$0.19
89.62
$\pm$0.24
72.71
$\pm$0.23
77.85
$\pm$1.03
54.69
$\pm$0.15
Block+EMoE 81.07
$\pm$0.53
46.33
$\pm$1.40
53.50
$\pm$0.97
79.62
$\pm$0.45
80.15
$\pm$0.61
79.03
$\pm$0.25
88.81
$\pm$0.19
91.33
$\pm$0.23
84.92
$\pm$0.07
88.87
$\pm$0.26
70.59
$\pm$1.00
74.62
$\pm$2.70
53.04
$\pm$1.82
EMoE 80.82
$\pm$0.10
44.18
$\pm$0.63
61.30
$\pm$2.24
76.93
$\pm$2.52
81.19
$\pm$0.27
77.87
$\pm$0.19
90.25
$\pm$0.10
92.14
$\pm$0.25
85.63
$\pm$0.30
90.21
$\pm$0.14
71.24
$\pm$1.22
78.00
$\pm$0.87
51.55
$\pm$0.88
EMoE+ln 81.04
$\pm$0.10
45.05
$\pm$0.88
58.17
$\pm$1.81
77.92
$\pm$1.64
81.27
$\pm$0.17
77.40
$\pm$0.25
90.08
$\pm$0.07
91.64
$\pm$0.54
85.67
$\pm$0.29
89.23
$\pm$0.67
72.14
$\pm$0.76
77.34
$\pm$1.33
52.81
$\pm$0.48
EMoE-learn 81.08
$\pm$0.24
46.09
$\pm$2.12
57.56
$\pm$0.43
77.52
$\pm$2.15
81.25
$\pm$0.13
78.16
$\pm$0.30
89.84
$\pm$0.13
91.87
$\pm$0.28
84.97
$\pm$0.33
89.76
$\pm$0.33
71.32
$\pm$2.03
77.47
$\pm$0.76
53.29
$\pm$0.45
EMoE-learn+ln
81.34
$\pm$0.30
44.46
$\pm$1.67
58.38
$\pm$0.59
77.78
$\pm$1.71
81.13
$\pm$0.26
78.28
$\pm$0.13
90.02
$\pm$0.29
91.48
$\pm$0.30
85.62
$\pm$0.46
89.87
$\pm$0.30
71.90
$\pm$0.61
77.38
$\pm$1.77
53.44
$\pm$0.47
BERT-Large (340 Million Parameters) Results (best result)
LoRA 80.75 45.42 58.24 82.03 81.79 78.46 90.35 92.05 85.95 90.37 72.06 78.41 54.44
LoRA+block 81.19 43.56 55.48 81.45 81.83 78.63 89.49 92.48 85.25 89.17 72.79 79.88 54.71
GMoE 81.78 49.14 59.72 79.55 81.66 78.41 90.31 91.86 85.93 90.42 72.06 79.02 55.16
Block+EMoE-learn
81.72 44.45 61.70 78.47 81.38 78.46 90.59 92.04 85.93 89.95 73.04 78.89 54.91
Block+EMoE 81.49 47.77 54.57 80.15 80.85 79.32 88.98 91.66 85.00 89.23 71.81 77.14 55.56
EMoE 80.95 45.05 64.31 79.64 81.55 78.10 90.37 92.42 86.03 90.34 72.55 79.23 52.24
EMoE+ln 81.17 45.71 60.73 79.84 81.48 77.63 90.15 92.33 86.01 90.60 72.79 79.53 53.28
EMoE-learn 81.34 48.75 58.11 79.89 81.25 78.58 89.93 92.23 85.38 90.21 74.02 78.33 53.91
EMoE-learn+ln
81.72 46.82 59.09 79.60 81.41 78.47 90.27 91.72 86.11 90.17 72.55 78.67 54.10
GPT2-XL (1.5 Billion Parameters) Results (mean and standard deviation)
LoRA 75.42
$\pm$2.71
41.78
$\pm$1.62
60.37
$\pm$1.32
77.36
$\pm$0.73
78.48
$\pm$0.33
78.57
$\pm$0.59
89.91
$\pm$0.73
90.05
$\pm$0.44
85.57
$\pm$1.34
88.85
$\pm$0.38
68.22
$\pm$1.27
73.60
$\pm$2.87
57.39
$\pm$0.34
LoRA+block 76.87
$\pm$2.47
41.09
$\pm$1.80
58.67
$\pm$0.97
78.57
$\pm$1.23
77.66
$\pm$0.25
78.78
$\pm$0.55
90.11
$\pm$0.71
91.54
$\pm$0.94
83.54
$\pm$0.57
88.85
$\pm$1.01
69.53
$\pm$0.64
68.22
$\pm$7.84
58.53
$\pm$0.46
GMoE 75.28
$\pm$3.67
42.50
$\pm$1.37
62.59
$\pm$0.37
77.89
$\pm$0.56
78.33
$\pm$0.58
79.18
$\pm$0.04
90.11
$\pm$0.38
88.64
$\pm$4.52
83.23
$\pm$0.40
89.47
$\pm$0.26
70.10
$\pm$1.39
74.61
$\pm$3.71
58.29
$\pm$0.78
Block+EMoE-learn
74.80
$\pm$5.15
44.15
$\pm$1.95
61.67
$\pm$1.18
78.18
$pm$1.04
78.27
$\pm$0.21
78.97
$\pm$0.20
90.22
$\pm$0.59
91.04
$\pm$0.65
83.82
$\pm$0.54
89.49
$\pm$0.67
69.93
$\pm$0.50
73.41
$\pm$3.94
57.58
$\pm$0.57
Block+EMoE 75.37
$\pm$3.04
39.33
$\pm$2.21
58.32
$\pm$2.40
72.05
$\pm$3.04
77.70
$\pm$0.59
79.32
$\pm$0.06
90.24
$\pm$0.36
91.79
$\pm$0.27
83.79
$\pm$0.27
89.23
$\pm$0.38
63.24
$\pm$8.56
70.60
$\pm$2.89
56.85
$\pm$0.57
EMoE 76.07
$\pm$2.12
42.95
$\pm$0.95
61.56
$\pm$1.65
78.09
$\pm$0.56
78.57
$\pm$0.67
78.87
$\pm$0.38
90.39
$\pm$0.55
91.87
$\pm$0.22
83.55
$\pm$1.04
89.14
$\pm$1.39
67.73
$\pm$1.55
74.06
$\pm$4.79
57.33
$\pm$0.59
EMoE+ln 74.35
$\pm$3.50
42.24
$\pm$0.89
61.88
$\pm$2.45
78.32
$\pm$0.95
78.67
$\pm$0.62
79.02
$\pm$0.30
89.51
$\pm$0.90
91.19
$\pm$0.56
83.12
$\pm$1.90
89.29
$\pm$1.09
69.20
$\pm$1.03
72.74
$\pm$5.10
57.62
$\pm$0.02
EMoE-learn 74.61
$\pm$4.75
40.50
$\pm$0.46
59.50
$\pm$3.74
80.04
$\pm$0.46
78.27
$\pm$0.59
79.06
$\pm$0.50
90.59
$\pm$0.66
91.69
$\pm$0.59
83.23
$\pm$1.54
89.79
$\pm$1.35
67.48
$\pm$1.30
73.62
$\pm$1.12
56.22
$\pm$0.15
EMoE-learn+ln
77.17
$\pm$1.58
42.52
$\pm$0.41
58.90
$\pm$4.16
79.56
$\pm$0.81
78.27
$\pm$0.54
79.19
$\pm$0.32
90.23
$\pm$0.86
91.98
$\pm$0.35
83.31
$\pm$1.08
88.50
$\pm$1.92
68.63
$\pm$1.51
71.26
$\pm$3.49
56.69
$\pm$0.37
GPT2-XL (1.5 Billion Parameters) Results (best result)
LoRA 78.17 43.93 61.77 78.39 78.87 79.12 90.52 91.03 84.81 89.39 69.12 77.48 57.79
LoRA+block 80.17 43.09 59.52 79.45 78.33 79.42 90.94 92.87 84.12 90.28 70.10 74.55 57.64
GMoE 80.23 43.96 63.05 78.68 78.00 79.33 90.57 92.27 83.69 89.81 71.08 79.79 59.36
Block+EMoE-learn
80.04 46.01 62.74 79.65 79.06 79.24 90.67 91.94 84.45 90.42 70.59 78.78 58.28
Block+EMoE 79.02 41.55 61.64 76.30 78.33 79.40 90.68 92.00 84.12 89.60 73.31 73.31 57.64
EMoE 77.58 43.84 63.78 78.75 79.29 79.26 91.17 92.19 84.42 90.22 69.85 80.81 58.01
EMoE+ln 77.92 43.39 65.32 79.16 79.25 79.35 90.48 91.74 85.73 90.79 70.59 79.81 57.87
EMoE-learn 79.35 41.15 64.56 80.40 78.76 79.76 91.32 92.47 84.49 90.77 69.12 74.77 56.42
EMoE-learn+ln
79.35 43.01 64.32 80.57 79.02 79.63 91.40 92.41 84.84 90.57 70.34 74.77 57.13
Table 14: Results of different MoEs Configurations
Algorithm MRPC CoLA RTE STSB Avg
Dense 86.83
$\pm$0.87
60.88
$\pm$2.54
78.70
$\pm$0.59
89.07
$\pm$0.11
78.87
EMoE 88.05
$\pm$0.35
63.11
$\pm$0.21
80.02
$\pm$0.34
89.37
$\pm$0.24
80.14
EMoE-learn 87.93
$\pm$0.61
62.87
$\pm$0.71
79.90
$\pm$0.61
89.40
$\pm$0.08
80.03
EMoE-last-every2
87.27
$\pm$0.47
61.60
$\pm$0.63
79.18
$\pm$0.17
89.38
$\pm$0.24
79.36
EMoE-learn-learn-every2
87.26
$\pm$0.21
61.82
$\pm$1.10
78.46
$\pm$1.22
89.31
$\pm$0.15
79.21
EMoE-every2
86.78
$\pm$0.34
59.21
$\pm$0.79
77.38
$\pm$1.04
89.31
$\pm$0.06
78.17
EMoE-learn-every2
86.71
$\pm$1.32
54.02
$\pm$0.47
74.25
$\pm$0.74
88.51
$\pm$0.31
75.87
Table 15: Comparison of EMoE and Dense Results with Different Training Data Fraction
Data CoLA MRPC RTE STSB SST2 QNLI Average
Fraction EMoE Dense EMoE Dense EMoE Dense EMoE Dense EMoE Dense EMoE Dense Diff
1.0 62.27 60.88 87.75 86.83 80.02 78.70 89.37 89.07 95.41 95.18 92.10 91.84 0.74
0.9 61.58 60.01 87.52 86.49 79.87 77.85 89.18 89.07 95.41 95.16 92.05 91.94 0.85
0.8 60.89 59.28 86.58 86.35 79.22 76.77 88.98 88.99 95.34 95.06 91.94 91.83 0.78
0.7 59.29 58.25 86.10 85.56 77.98 76.77 87.95 87.95 95.41 95.06 92.04 91.74 0.57
0.6 58.91 57.93 85.61 84.76 76.29 75.45 86.70 86.17 95.26 95.03 91.83 91.54 0.62
0.5 55.18 53.89 84.91 84.76 76.53 75.21 85.63 85.59 95.19 94.91 91.23 91.12 0.53
0.3 50.17 50.29 82.80 82.59 73.52 72.44 80.23 79.08 94.82 94.72 90.45 90.26 0.44
0.1 46.47 45.54 78.17 77.85 63.05 63.17 62.33 60.81 94.49 94.38 88.39 88.32 0.47
A.5 More Visualization Results
In this part, we demonstrate more gating visualization results on SST-2, STSB, MRPC, and RTE in Figure 8. These results are consistent with earlier findings: (1) Both $\operatorname{avg-k}$ gating and learned gating converge, as indicated by the lower halves of the plots. (2) $\operatorname{avg-k}$ gating is more stable than learned gating. This could mitigate data inefficiency resulting from inconsistencies in gating across different stages of training (Zuo et al., 2022).
Figure 8: Expert selections during training with distinct gating functions ($\operatorname{avg-k}$ vs learned gate) and expert types (modules from dense vs repetitions of dense). The vertical axis illustrates training steps, with top-down arrangement signifying begin-end; the horizontal axis represents expert selection frequency within 1K steps. (a), (b) and (c) correspond respectively to EMoE, EMoE-learn, and GMoE configurations. The subplots from top to bottom are results for SST-2, STS-B, MRPC, and RTE.